Suppose that the price of an ice cream cone increases from PRs. 60 to PRs. 80 and the amount you buy falls from 12 to 8 cones.

Requirements:

a. Calculate price elasticity of demand in Percentage

b. Interpret the answer

c. What type of elasticity is this i.e. perfectly elastic, perfectly inelastic, unitary elastic,

relatively elastic or relatively inelastic.


1
Expert's answer
2021-01-29T13:27:12-0500

a) By the definition of the price elasticity of demand, we have:


"E_d=\dfrac{\%\Delta Q}{\%\Delta P},""E_d=\dfrac{\dfrac{12-8}{12}\cdot100\%}{\dfrac{80-60}{60}\cdot100\%}=\dfrac{33.3\%}{33.3\%}=1."

b) As we can see from the calculations, the demand changes in the same proportion to its price (the percentage change in demand is exactly equal to the percentage change in price).

c) According to the interpretation above there is a unitary elasticity.